amy drove 200 miles at an average speed 10 miles per hour faster than her usual speed. if she completed the trip in 1 hour less time than usual, what is her usual driving speed in miles per hour?
Found 2 solutions by stanbon, mamiya:
Answer by stanbon(75887)

amy drove 200 miles at an average speed 10 miles per hour faster than her usual speed. if she completed the trip in 1 hour less time than usual, what is her usual driving speed in miles per hour?
---
Early trip DATA:
distance = 200 miles; rate = r+10 mph ; time = 200/(r+10) hrs
------------------------
Later trip DATA:
distance = 200 miles : rate = r mph ; time = 200/r hrs
-------------------------
Equation:
early time - later time = 1 hr
---
200/r - 200/(r+10) = 1 hr.
---
200(r+10) - 200r = r(r+10)
----
200*10 = r^2+10r
----
r^2 + 10r - 2000 = 0
----
(r-40)(r+50) = 0
======================
Positive Solution:
rate = 40 mph
======================
Cheers,
Stan H.
==========

Answer by mamiya(56)

let x be the usual average speed and t the among of times she usually took to complete the trip. let s be the average speed , t' the among of time she takes now and d the distance.
s= x + 10 since her average speed now is 10 miles per hour more than her usual speed
d= 200 miles
t'= t - 1 since she completed the trip in 1 hour less time usual
we know that distance= speed x time
so d= st' and d= xt
so xt= 200
(x+10)(t-1)=200
xt=200
xt - x + 10t -10= 200
xt=200
200 - x + 10t - 10 =200
xt=200
10t - x= 10

x= 10(t - 1)
xt=200
x= 10(t-1)
10(t-1)*t = 200
10t^2 - 10t - 200 =0

t^2 - t - 20 =0
let's call p(x)= t^2 - t - 20 and let's factorize this polynomial.
the determinant D= (-1)^2 - 4(1)(-20)= 81= (9)^2
t1= (1+9)/2 = 5
t2= (1-9)/2 = -4
so the answers of the equation ( t^2-T-20=0) are 5 and -4, but since t is an among of time , t has to be a positive number, so answer can't be -4 , therefore the answer is 5.
so t=5
x=10(t-1)
= 10(5-1)
= 40
so, her usual driving speed is 40 miles per hour.